It takes 7 men 10 days to build a garage. Half way through building the garage they get some help from 3 more men and continue to work at the same rate.
a) How long will it take to FINISH OFF the garage? __________________________________
b) How long did it take to build the garage altogether? ________________________________
I tried finding how many days 1 man takes to build the garage but I do not know how to do that...

Answer by ptaylor(2198)

Let x=amount of time it takes to finish off the garage
It takes 7*10=70 man-days to build the garage
Halfway through building the garage would take 5 days; 7*5=35 man-days
This would leave 35 man-days to complete by 7+3=10 men
Soooo 10x=35
x=3.5 days------ time to finish off the garage (a)
5+3.5=8.5 days----time to build the garage altogether (b)
BTW, one man takes 70 days to build the garage
Hope this helps---ptaylor